Driving down the road at a speed of 40.0 m/s, you suddenly notice a fallen tree blocking the road a distance of 88.4 m ahead of you. You step on the brake pedal and decelerate at a constant rate. What must the magnitude of your acceleration be so that you will come to a stop 5.9 m in front of the tree?
Answer #1

Total distance, d = 88.4 - 5.9 = 82.5 m

Using 3rd equation of motion,

v^2 = u^2 + 2as

0 = 40^2 - 2 x a x 82.5

Acceleration, a = 9.7 m/s^2

Comment in case any doubt please rate my answer....